A commonly asked question is, “How long will it take to double my money?” At 7% interest rate and continuous compounding, what is the answer? Round your answer to the hundredth's place.
Answer by stanbon(26297)

A commonly asked question is, “How long will it take to double my money?” At 7% interest rate and continuous compounding, what is the answer? Round your answer to the hundredth's place.
A(t) = Pe^(rt)
----
If P is doubled you get:
2P = Pe^(rt)
e^(rt) = 2
---
Take the natural log of both sides:
rt = ln(2)
t = [ln(2)]/r
t = [ln(2)]/0.07
t = 9.9 years (time to double the money at 7% interest compouded continuously)
====================================================
Cheers,
Stan H.